# Adversarial bandit optimization for approximately linear functions

### v5

@@ 0.A.3 Proof of Lemma 8
$\min_{a\in\mathcal{K}}\|x^{*}-a\|\leq\|x^{*}-(1-\delta)x^{*}\|\leq\delta D.$ (30)
@@
By Cauchy–Schwarz inequality and the fact that $\|\theta\|\leq G$ and $\|x_{\delta}^{*}-x^{*}\|\leq\delta D$,
@@
[…]
@@
$\leq$ $\sqrt{d^{2}\mathbb{E}_{t}[\sigma_{t}^{2}(y_{t})\mu_{t}^{\top}\mu_{t}]}$ (37)
@@
Then we bound $\|x_{\delta}^{*^{\prime}}-x_{t}^{\prime}\|_{x_{t}^{\prime}}$. From the triangle inequality,
$\|x_{\delta}^{*^{\prime}}-x_{t}^{\prime}\|_{x_{t}^{\prime}}$$\leq$ $\|x_{\delta}^{*^{\prime}}\|_{x_{t}^{\prime}}+\|x_{t}^{\prime}\|_{x_{t}^{\prime}}.$
By Assumption 1, Lemma 1 and Lemma 4, we obtain $\|x_{\delta}^{*^{\prime}}\|_{x_{t}^{\prime}}\leq 2\nu$, $\|x_{t}^{\prime}\|_{x_{t}^{\prime}}=\sqrt{\nu}$
So $\|x_{\delta}^{*^{\prime}}-x_{t}^{\prime}\|_{x_{t}^{\prime}}\leq 2\nu+\sqrt{\nu}$ and $\sum_{t=1}^{T}M_{t}^{\top}(x_{\delta}^{*^{\prime}}-x_{t}^{\prime})\leq Td\epsilon(2\nu+\sqrt{\nu})$. Then bound $\sum_{t=1}^{T}\mathbb{E}_{t}[g_{t}]^{T}(x_{t}^{\prime}-x_{\delta}^{*^{\prime}})$
@@
$\sum_{t=1}^{T}\theta_{t}^{\top}x_{t}-\sum_{t=1}^{T}\theta_{t}^{\top}x^{*}_{\delta}\leq 4\eta d^{2}T+\frac{\nu\log(\frac{1}{\delta})}{\eta}+Td\epsilon(2\nu+\sqrt{\nu}).$ (43)

### v8

@@ 0.A.3 Proof of Lemma 8
$\min_{a\in\mathcal{K_{\delta}}}\|x^{*}-a\|\leq\|x^{*}-(1-\delta)x^{*}\|\leq\delta D.$ (30)
@@
By Cauchy–Schwarz inequality and the fact that $\|\theta_{t}\|\leq G$ and $\|x_{\delta}^{*}-x^{*}\|\leq\delta D$,
@@
Let $\theta_{t}^{\prime}=(\theta_{t},z)$, where $z$ is the $(d+1)$th coordinate of $d\mathbb{E}_{t}[(\theta_{t},0)^{\top}(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}]$. From Lemma 6, we know $d\mathbb{E}_{t}[(\theta_{t},0)^{\top}(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}]=\theta_{t}^{\prime}$. Since $g_{t}=df(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}=d(\theta_{t},0)^{\top}(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}+d\sigma_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}$, and letting $M_{t}=\mathbb{E}_{t}[d\sigma_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}]$, it follows that $\theta_{t}^{\prime}=\mathbb{E}_{t}[g_{t}]-M_{t}$. Consequently,
@@
$=$ $\sqrt{d^{2}\|\mathbb{E}_{t}[\sigma_{t}(y_{t})\mu_{t}]\|^{2}}$ (37)
$\leq$ $\sqrt{d^{2}\mathbb{E}_{t}[\|\sigma_{t}^{2}(y_{t})\mu_{t}^{\top}\mu_{t}\|]}$ (38)
@@
Then we bound $\|x_{\delta}^{*^{\prime}}-x_{t}^{\prime}\|_{x_{t}^{\prime}}$. By Lemma 4,
$\|x_{\delta}^{*^{\prime}}-x_{t}^{\prime}\|_{x_{t}^{\prime}}$ $\leq$ $2(\nu+2\sqrt{\nu})\frac{(1-\delta)}{\delta}.$
So $\sum_{t=1}^{T}M_{t}^{\top}(x_{\delta}^{*^{\prime}}-x_{t}^{\prime})\leq Td\epsilon 2(\nu+2\sqrt{\nu})\frac{(1-\delta)}{\delta}$. Then bound $\sum_{t=1}^{T}\mathbb{E}_{t}[g_{t}]^{T}(x_{t}^{\prime}-x_{\delta}^{*^{\prime}})$
@@
$\sum_{t=1}^{T}\theta_{t}^{\top}x_{t}-\sum_{t=1}^{T}\theta_{t}^{\top}x^{*}_{\delta}\leq 4\eta d^{2}T+\frac{\nu\log(\frac{1}{\delta})}{\eta}+Td\epsilon 2(\nu+2\sqrt{\nu})\frac{(1-\delta)}{\delta}+\delta DGT.$ (44)